OpenStudy (anonymous):

2 questions in calculus:

OpenStudy (abb0t):

Why not one at a time?

OpenStudy (anonymous):

prove: 1) \[\lim_{x \rightarrow 0+} \frac{x}{x-\left[ sinx \right]} = 1\] and \[\lim_{x \rightarrow 0-} \frac{x}{x-\left[ sinx \right]} = 0\] 2) Prove: \[\lim_{x \rightarrow \frac{ \Pi }{2}}\frac{1}{cosx} \neq \infty\]

OpenStudy (anonymous):

my bad abb0t, I believe these are 2 easy ones... but I'll ask 1 at a time from now on

OpenStudy (abb0t):

Multiply the first one by the conjugate

OpenStudy (abb0t):

Also, \(\frac{ 1 }{ \cos(x) } = \sec(x)\)

OpenStudy (anonymous):

we're supposed to prove using the epsilon/delta definition of limit

OpenStudy (anonymous):

and I can't seem to finish it... got question 1 to this: \[|\frac{sinx}{x-sinx}|\] and now I need to define Delta somehow

OpenStudy (abb0t):

For every \(\epsilon >0\) there exists some \(\delta >0\) such that \(|f(x)-L|< \epsilon \) whenever \(0<|x-a|< \delta\)

OpenStudy (abb0t):

L in the first case is 1

OpenStudy (anonymous):

and in question 2 I can't find how to define M so \[\frac{1}{cosx}\le N\]

OpenStudy (anonymous):

i know how to define but I have a problem with the steps you take to figure out how to pick the specific delta/epsilon

OpenStudy (abb0t):

\(\left| \frac{ x }{ x-\sin(x) } - 1 \right| < \epsilon\) whenever \(0<|x-0|< \delta \) do the same with zero now. Use algebra to solve for epsilon basically.

OpenStudy (abb0t):

since epsilon is just some arbitrary number.

OpenStudy (anonymous):

I did this and got to \[|\frac{sinx}{x-sinx}|\] when I know that |x|<delta and now i'm stuck on how to define delta so this expression will be smaller than epsilon

OpenStudy (anonymous):

I usually get stuck on changing the expresion using algebra

OpenStudy (anonymous):

same with the other question, I know what I'm supposed to do but I'm stuck on how to 'play" with equation